OpenStudy (jokersmart):

Check my answer? What is the value of x? 3x^3-6=186 4?

OpenStudy (jokersmart):

Equation is properly wrote like this: \[3x^{3}-6=18\]

OpenStudy (jokersmart):

I'm sorry it is 186* at the end.

OpenStudy (valpey):

Add 6 to both sides. Divide both sides by 3...

OpenStudy (johnweldon1993):

Ahh 186 that makes sense... \[\large 3x^3 - 6 = 186\] \[\large 3x^3 = 192\] \[\large x^3 = 64\] \[\large x = \sqrt[3]{64} = 4\] Correct!

OpenStudy (jokersmart):

Thank you @johnweldon1993 ! :)